### v4

@@ 3.1.2 Face Reenactment
This section reviews current methods from four points: 3DMM-based, landmark matching, face feature decoupling, and self-supervised learning. We summarize them in Table 3.
@@
$\bullet$ 3DMM-based. Some methods koujan2020head2head ; yang2022face2face ; gao2023high utilize 3DMM to construct a facial parameter model as an intermediary for transferring information between the source and target. In particular, Face2Faceρ yang2022face2face , based on 3DMM, consists of a u-shaped rendering network driven by head pose and facial motion fields and a hierarchical coarse-to-fine motion network guided by landmarks at different scales. However, some methods thies2016face2face ; kim2018deep ; kim2019neural exhibit visible artifacts in the background when dealing with significant head movements in the input images. To address issues such as incomplete attribute decoupling in facial reproduction tasks, PECHead gao2023high models facial expressions and pose movements. It combines self-supervised learning of landmarks with 3D facial landmarks and introduces a new motion-aware multi-scale feature alignment module to eliminate artifacts that may arise from facial motion.
@@
$\bullet$ Landmark Matching. This kind of methods wiles2018x2face ; zakharov2020fast ; ha2020marionette ; hsu2022dual ; doukas2023free aim to establish a mapping relationship between semantic objects in the facial regions of the driving source and the target source through landmarks. Based on this mapping relationship zakharov2019few ; zhang2020freenet , the transfer of facial movement information is achieved. In particular, X2Face wiles2018x2face achieves one-to-one driving of facial expressions and poses for each frame of the character in the driving video and source video. To address the challenge of reproducing large head poses in facial reenactment, Xu et alhsu2022dual propose a dual-generator network incorporating a 3D landmark detector into the model. Free-headgan doukas2023free comprises a 3D keypoint estimator, an eye gaze estimator, and a generator built on the HeadGAN architecture. The 3D keypoint estimator addresses the regression of deformations related to 3D poses and expressions. The eye gaze estimator controls eye movement in videos, providing finer details. MetaPortrait zhang2023metaportrait achieves accurate distortion field prediction through dense facial keypoint matching and accelerates model training based on meta-learning principles, delivering excellent results on limited datasets.
$\bullet$ Feature Decoupling. The latent feature decoupling and driving methods bounareli2023hyperreenact ; bounareli2023stylemask ; li2023one ; bounareli2024one ; yang2023learning aims to disentangle facial features in the latent space of the driving video, replacing or mapping the corresponding latent information to achieve high-fidelity facial reproduction under specific conditions. HyperReenact bounareli2023hyperreenact uses attribute decoupling, employing a hyper-network to refine source identity features and modify facial poses. StyleMask bounareli2023stylemask separates facial pose and expression from the identity information of the source image by learning masks and blending corresponding channels in the pre-trained style space S of StyleGAN2. HiDe-NeRF li2023one employs a deformable neural radiance field to represent a 3D scene, with a lightweight deformation module explicitly decoupling facial pose and expression attributes.
$\bullet$ Self-supervised Learning. Self-supervised learning employs supervisory signals inferred from the intrinsic structure of the data, reducing the reliance on external data labels tripathy2020icface ; zeng2020realistic ; oorloff2023robust ; zhang2023exploiting . Oorloff et aloorloff2023robust employs self-supervised methods to train an encoder, disentangling identity and facial attribute information of portrait images within the pre-defined latent space itself of a pre-trained StyleGAN2. Zhang et alzhang2023exploiting utilizes 3DMM to provide geometric guidance, employs pre-computed optical flow to guide motion field estimation, and relies on pre-computed occlusion maps to guide the perception and repair of occluded areas.
@@
Table 4: Overview of representative talking face generation methods. Notations: ➊ LRW, ➋ VoxCeleb2, ➌ MEAD, ➍ Self-build, ➎ LRS2, ➏ HDTF, ➐ LRS3, ➑ CREMA-D, ➒ VoxCeleb, ➓ FFHQ.
@@
Audio / Text - Driven Chen et alchen2018lip ECCV’18 cooke2006audio 0Audiovisual Poor resolution, inability to control pose and emotional. Proposed a novel generator network and a comprehensive model with four complementary losses, as well as a new audio-visual related loss function to guide video generation.
@@
Zhou et alzhou2019talking AAAI’19 Inability to control pose and emotional variations. Generate high-quality talking face videos by disentangling audio-visual representations.
@@
Chen et alchen2019hierarchical CVPR’19 Inability to control pose and emotional variations. Proposed a cascaded approach, using facial landmarks as an intermediate high-level representation.
@@
Wav2Lip prajwal2020lip ICMR’20 ➊➎➐ Poor resolution, inability to control pose and emotional. A new evaluation framework and a dataset for training mouth synchronization are proposed.
@@
MakeItTalk zhou2020makelttalk TOG’20 Uable to control pose and emotional variations well. Separating content information and identity information from audio signals, combining LSTM and self -attention mechanism to enhance head movement coherence.
@@
Ji et alji2021audio CVPR’21 ➊➌ Inability to control pose and emotional variations. By breaking down the input audio sample into content and emotion embeddings, cross-reconstruction of emotional disentanglement creates facial landmarks with nuanced emotional content.
@@
SPACE gururani2023space ICCV’23 ➋➌ Lack emotional and other latent attributes control Constructed a novel facial intermediate representation, achieving control overhead pose, blinking, and gaze direction.
Gan et algan2023efficient ICCV’23 ➊➋➌ Insufficient emotional output diversity. Proposed a two-stage architecture, implementing speaker generation independent of emotion and emb -edding emotion information.
@@
SadTalker zhang2023sadtalker CVPR’23 ➏➒ Lack emotional and other latent attributes control. Based on the idea of 3DMM and conditional VAE, 3D coefficients controlling facial motion and expre -ssion are generated from audio to realize the reproduction of accurate faces from audio.
Zhongetczhong2023identity CVPR’23 ➎➐ afouras2018deep Lack emotional and other latent attributes control. A two-stage framework for step-by-step audio-to-face action generation is proposed.
@@
EmoTalk peng2023emotalk ICCV’23 ➍➏ Poor real-time performance and expression details. An emotion-entangled encoder and emotion-guided decoder enable emotion injection, with outputs generated using Blendshape and FLAME model rendering.
@@
TalkLip wang2023seeing CVPR’23 ➊➎ Inability to control pose and emotional variations. Pre-trained lip-reading experts are employed to penalize incorrect lip-reading predictions in the synth -esized videos.
@@
DR2 zhang2024dr2 WACV’24 Lack emotional and other latent attributes control. The model explored effective strategies for reducing the training workload.
@@
RADIO lee2024radio WACV’24 ➊➋➏ Lack emotional and other latent attributes control. Introducing StyleGAN2 style modulation to adapt to human identity and utilizes ViT blocks to focus on facial attributes in the reference image.
@@
Multimodal PC-AVS zhou2021pose CVPR’21 ➊➋ Lack emotional and other latent attributes control. Introduction of pose-source video drive compensation to generate head motion in video.
@@
GC-AVT liang2022expressive CVPR’22 ➋➌ Poor resolution, unable to handl complex backgrounds. In addition to the source image, a gesture source, an expression source, and audio are introduced to drive the talking head generation.
@@
Yu et alyu2021multimodal TMM’22 Lack emotional and other latent attributes control. Fusion of audio and text inputs for more accurate lip movement and chin posture prediction.
@@
Xu et alxu2023high CVPR’23 Insufficient control over the intensity of emotional output. Embedding textual, visual, and auditory emotional modalities into a unified space.
@@
LipFormer wang2023lipformer CVPR’23 ➎➓ Poor ability to preserve face feature attributes. Propose retaining high-quality facial details obtained from pre-training in a codebook format and repro -ducing them by driving the encoded mapping relationship between audio and lip movements.
@@
Wang et alwang2024styletalk++ TPAMI’24 ➋➌➏ Unable to delicately control emotions. Using 3DMM as an intermediate variable to convey facial expressions and head movements, and intro -ducing additional reference videos to extract the desired speaking style.
@@
Diffusion DAE-Talker du2023dae MM’23 High model complexity. It replaces traditional manually crafted intermediate representations with data-driven latent representa -tions obtained from a DAE.
@@
Yu et alyu2023talking ICCV’23 ➋➒ Poor resolution, high model complexity. Building a corresponding mapping between audio and non-lip representations and training using the diffusion model.
DreamTalk ma2023dreamtalk arXiv’23 ➋➌➏ Mismatched emotion and semantics, occasional artifacts. The denoising network, style-aware lip expert, and style predictor collaborate to make the model perform well in various speaking styles.
@@
Stypułkowski stypulkowski2024diffused WACV’24 ➊➑ High model complexity, short video generation duration. The model incorporates motion frame and audio embedding information to capture past movements and future expressions, with an emphasis on the mouth region through an additional lip sync loss.
@@
EmoTalker zhang2024emotalker ICASSP’24 ➌➑ High model complexity. It achieves emotion-editable talking face generation based on a conditional diffusion model.
@@
VASA-1 xu2024vasa arXiv’24 High model complexity. Expressive and well-decoupled facial latent space has been constructed, and highly controllable, high -quality generation effects have been achieved based on the Diffusion Transformer.
@@
3D-Model AD-NeRF guo2021ad ICCV’21 Inadequate control of emotions and latent attributes. The NeRF based approach achieves accurate reproduction of detailed facial components and generates the upper body region.
@@
DFRF shen2022learning ECCV’22 Lack of emotional and other latent attributes control. Combining audio with 3D perceptual features and proposing an facial deformation module.
@@
AE-NeRF li2023ae AAAI’24 Lack emotional and other latent attributes control. Facial modeling is divided into NeRF related to audio and unrelated to audio to enhance audio-visual lip synchronization and facial detail.
@@
SyncTalk peng2023synctalk CVPR’24 Lack controllable emotional intensity regulation. The facial sync controller boosts component coordination, and a portrait generator corrects artifacts, enhancing video details.
@@
Ye et alye2024real3d ICLR’24 zhu2022celebv Lack emotional control and occasional artifacts. Facial and audio information is separately represented using tri-plane, followed by rendering. The gen -erated results are further optimized based on the super-resolution network.
3.1.3 Talking Face Generation
@@
In this section, we review current methods from three perspectives: audio/text driven, multimodal conditioned, diffusion-based, and 3D-model Technologies. We also summarize them in Table 4.
@@
$\bullet$ Audio/Text Driven. Methods aim to map and guide lip and facial movements in generated videos by understanding the semantic information from the driving source apb2face ; shen2022learning ; wang2023seeing . Early methods fan2015photo ; chen2018lip perform poorly in terms of generalization and training complexity. After training, the models struggled to generalize to new individuals, requiring extensive conversational data for training new characters. Researchers chen2019hierarchical ; prajwal2020lip propose their solutions from various perspectives. However, Most of these methods prioritize generating lip movements aligned with semantic information, overlooking essential aspects like identity and style, such as head pose changes and movement control, which are crucial in natural videos. To address this, MakeItTalk zhou2020makelttalk decouples input audio information by predicting facial landmarks based on audio and obtaining semantic details on facial expressions and poses from audio signals. SadTalker zhang2023sadtalker extracts 3D motion coefficients for constructing a 3DMM from audio and uses this to modulate a new 3D perceptual facial rendering for generating head poses in talking videos. Additionally, some methods wu2023speech2lip ; wang2023memory ; lee2024radio ; fu2023mimic propose their improvement methods, and these will not be detailed one by one. In addition, the emotional expression varies for different texts during a conversation, and vivid emotions are an essential part of real talking face videos sheng2023stochastic ; gan2023efficient . Recently, some methods tan2023emmn ; zhai2023talking ; gururani2023space extend their previous approaches by incorporating matching between the driving information and corresponding emotions. EMMN tan2023emmn establishes an organic relationship between emotions and lip movements by extracting emotion embeddings from the audio signal, synthesizing overall facial expressions in talking faces rather than focusing solely on audio for facial expression synthesis. AMIGO zhai2023talking employs a sequence-to-sequence cross-modal emotion landmark generation network to generate vivid landmarks aided by audio information, ensuring that lips and emotions in the output image sequence are synchronized with the input audio. However, existing methods still lack effective control over the intensity of emotions. In addition, TalkCLIP ma2023talkclip introduces style parameters, expanding the style categories for text-guided talking video generation. Zhong et alzhong2023identity propose a two-stage framework, incorporating appearance priors during the generation process to enhance the model’s ability to preserve attributes of the target face. DR2 zhang2024dr2 explores practical strategies for reducing the training workload.
@@
$\bullet$ Multimodal Conditioned. To generate more realistic talking videos, some methods zhou2021pose ; liang2022expressive ; xu2023high ; wang2023lipformer introduce additional modal information on top of audio-driven methods to guide facial pose and expression. GC-AVT liang2022expressive generates realistic talking videos by independently controlling head pose, audio information, and facial expressions. This approach introduces an expression source video, providing emotional information during the speech and the pose source video. However, the video quality falls below expectations, and it struggles to handle complex background changes. Xu et alxu2023high integrate text, image, and audio-emotional modalities into a unified space to complement emotional content in textual information. Multimodal approaches have significantly enhanced the vividness of generated videos, but there is still room for exploration of organically combining information driven by different sources and modalities.
@@
$\bullet$ Diffusion-based. Recently, some methods mir2023dit ; ma2023dreamtalk ; du2023dae ; xu2023multimodal ; zhang2024emotalker apply the Diffusion model to the task of talking face generation. For fine-grained talking video generation, DAE-Talker du2023dae replaces manually crafted intermediate representations, such as facial landmarks and 3DMM coefficients, with data-driven latent representations obtained from a Diffusion Autoencoder (DAE). The image decoder generates video frames based on predicted latent variables. EmoTalker zhang2024emotalker utilizes a conditional diffusion model for emotion-editable talking face generation. It introduces emotion intensity blocks and the FED dataset to enhance the model’s understanding of complex emotions. Very recently, diffusion models are gaining prominence in talking face generation tasks stypulkowski2024diffused ; wei2024aniportrait ; tian2024emo and video generation tasks wang2023leo ; karras2023dreampose . Emo tian2024emo directly predicts video from audio without the need for intermediate 3D components, achieving excellent results. However, the lack of explicit control signals may easily lead to unnecessary artifacts. Based on the Diffusion Transformer architecture, VASA-1 xu2024vasa finely encodes and reconstructs facial details, constructing an expressive and well-decoupled facial latent space. This results in highly controllable and high-resolution video generation.
@@
$\bullet$ 3D-model Technologies. 3D models, exemplified by NeRF, are gaining traction in talking face generation guo2021ad ; shen2022learning ; li2023ae . AD-NeRF guo2021ad directly feeds features from the input audio signal into a conditional implicit function to generate a dynamic NeRF. AE-NeRF li2023ae employs a dual NeRF framework to separately model audio-related regions and audio-independent regions. Furthermore, some methods peng2023synctalk ; ye2024real3d adopt Tri-Plane chan2022efficient to represent facial and audio attributes. Synctalk peng2023synctalk models and renders head motion using a tri-plane hash representation, and then further enhances the output quality using a portrait synchronization generator. Very recently, 3D Gaussian Splatting kerbl20233d also been widely applied to this task. Some method chen2024gstalker ; cho2024gaussiantalker ; li2024talkinggaussian ; yu2024gaussiantalker introduce 3DGS to achieve more refined facial reconstruction and motion details, aiming to address the issue of insufficient pose and expression control caused by NeRF’s implicit representation.
@@
Table 5: Overview of representative facial attribute editing methods. Notations: ➊ FFHQ, ➋ CelebA, ➌ CelebA-HQ, ➍ CelebAMask-HQ, ➎ VoxCeleb, ➏ CelebAText-HQ, ➐ LFW, ➑ MM CelebA-HQ, ➒ CARLA, ➓ Multi-PIE. In addition, some abbreviations are used in the table: SIGGRAPH (SIG.), GANs (G.), Diffusion (D.), Transformer (T.).
@@
SC-FEGAN jo2019sc ICCV’19 G. Users can generate high-quality edited output images by freely sketching parts of the source image.
@@
AttGAN he2019attgan TIP’19 G. Applying attribute classification constraints to generated images has validated the drawbacks of enforcing stringent attribute independence constraints in latent representations.
@@
Shen et alshen2020interpreting CVPR’20 G. Thoroughly investigated how to encode different semantics in the latent space and explored the disentanglement between various semantics to achieve precise control over facial attributes.

### v5

@@ 3.1.2. Face Reenactment
This section reviews current methods from four points: 3DMM-based, landmark matching, face feature decoupling, and self-supervised learning. We summarize them in Tab. 3.
@@
$\bullet$ 3DMM-based. Some methods (yang2022face2face, ; gao2023high, ) utilize 3DMM to construct a facial parameter model as an intermediary for transferring information between the source and target. In particular, Face2Faceρ (yang2022face2face, ), based on 3DMM, consists of a u-shaped rendering network driven by head pose and facial motion fields and a hierarchical coarse-to-fine motion network guided by landmarks at different scales. However, some methods (thies2016face2face, ; kim2018deep, ; kim2019neural, ) exhibit visible artifacts in the background when dealing with significant head movements in the input images. To address issues such as incomplete attribute decoupling in facial reproduction tasks, PECHead (gao2023high, ) models facial expressions and pose movements. It combines self-supervised learning of landmarks with 3D facial landmarks and introduces a new motion-aware multi-scale feature alignment module to eliminate artifacts that may arise from facial motion.
@@
$\bullet$ Landmark Matching. This kind of methods (wiles2018x2face, ; zakharov2020fast, ; ha2020marionette, ; hsu2022dual, ; doukas2023free, ) aim to establish a mapping relationship between semantic objects in the facial regions of the driving source and the target source through landmarks. Based on this mapping relationship (zakharov2019few, ; zhang2020freenet, ), the transfer of facial movement information is achieved. To address the challenge of reproducing large head poses in facial reenactment, Xu et al(hsu2022dual, ) propose a dual-generator network incorporating a 3D landmark detector into the model. Free-headgan (doukas2023free, ) comprises a 3D keypoint estimator, an eye gaze estimator, and a generator built on the HeadGAN architecture. The 3D keypoint estimator addresses the regression of deformations related to 3D poses and expressions. The eye gaze estimator controls eye movement in videos, providing finer details. MetaPortrait (zhang2023metaportrait, ) achieves accurate distortion field prediction through dense facial keypoint matching and accelerates model training based on meta-learning principles, delivering excellent results on limited datasets.
@@
Table 4. Overview of representative talking face generation methods. Notations: ➊ LRW, ➋ VoxCeleb2, ➌ MEAD, ➍ Self-build, ➎ LRS2, ➏ HDTF, ➐ LRS3, ➑ CREMA-D, ➒ VoxCeleb, ➓ FFHQ.
@@
Audio / Text - Driven Chen et al(chen2018lip, ) ECCV’18 (cooke2006audio, )(0Audiovisual, ) Poor resolution, inability to control pose and emotional. Proposed a novel generator network and a comprehensive model with four complementary losses, as well as a new audio-visual related loss function to guide video generation.
@@
Zhou et al(zhou2019talking, ) AAAI’19 Inability to control pose and emotional variations. Generate high-quality talking face videos by disentangling audio-visual representations.
@@
Chen et al(chen2019hierarchical, ) CVPR’19 Inability to control pose and emotional variations. Proposed a cascaded approach, using facial landmarks as an intermediate high-level representation.
@@
Wav2Lip (prajwal2020lip, ) ICMR’20 ➊➎➐ Poor resolution, inability to control pose and emotional. A new evaluation framework and a dataset for training mouth synchronization are proposed.
@@
MakeItTalk (zhou2020makelttalk, ) TOG’20 Uable to control pose and emotional variations well. Separating content information and identity information from audio signals, combining LSTM and self -attention mechanism to enhance head movement coherence.
@@
Ji et al(ji2021audio, ) CVPR’21 ➊➌ Inability to control pose and emotional variations. By breaking down the input audio sample into content and emotion embeddings, cross-reconstruction of emotional disentanglement creates facial landmarks with nuanced emotional content.
@@
SPACE (gururani2023space, ) ICCV’23 ➋➌ Lack emotional and other latent attributes control Constructed a novel facial intermediate representation, achieving control overhead pose, blinking, and gaze direction.
@@
SadTalker (zhang2023sadtalker, ) CVPR’23 ➏➒ Lack emotional and other latent attributes control. Based on the idea of 3DMM and conditional VAE, 3D coefficients controlling facial motion and expre -ssion are generated from audio to realize the reproduction of accurate faces from audio.
@@
EmoTalk (peng2023emotalk, ) ICCV’23 ➍➏ Poor real-time performance and expression details. An emotion-entangled encoder and emotion-guided decoder enable emotion injection, with outputs generated using Blendshape and FLAME model rendering.
@@
TalkLip (wang2023seeing, ) CVPR’23 ➊➎ Inability to control pose and emotional variations. Pre-trained lip-reading experts are employed to penalize incorrect lip-reading predictions in the synth -esized videos.
@@
DR2 (zhang2024dr2, ) WACV’24 Lack emotional and other latent attributes control. The model explored effective strategies for reducing the training workload.
@@
RADIO (lee2024radio, ) WACV’24 ➊➋➏ Lack emotional and other latent attributes control. Introducing StyleGAN2 style modulation to adapt to human identity and utilizes ViT blocks to focus on facial attributes in the reference image.
FT2TF (diao2025ft2tf, ) WACV’25 ➎➐ Lack controllable emotional intensity regulation. A one-stage pipeline that generates realistic talking faces by integrating visual and textual input.
Talkclip (ma2025talkclip, ) TMM’25 ➋➌➏ Insufficient control over the intensity of emotional output. The method can use text to modulate expression intensity and edit expressions.
@@
Multimodal PC-AVS (zhou2021pose, ) CVPR’21 ➊➋ Lack emotional and other latent attributes control. Introduction of pose-source video drive compensation to generate head motion in video.
@@
GC-AVT (liang2022expressive, ) CVPR’22 ➋➌ Poor resolution, unable to handl complex backgrounds. In addition to the source image, a gesture source, an expression source, and audio are introduced to drive the talking head generation.
@@
Yu et al(yu2021multimodal, ) TMM’22 Lack emotional and other latent attributes control. Fusion of audio and text inputs for more accurate lip movement and chin posture prediction.
@@
Xu et al(xu2023high, ) CVPR’23 Insufficient control over the intensity of emotional output. Embedding textual, visual, and auditory emotional modalities into a unified space.
@@
LipFormer (wang2023lipformer, ) CVPR’23 ➎➓ Poor ability to preserve face feature attributes. Propose retaining high-quality facial details obtained from pre-training in a codebook format and repro -ducing them by driving the encoded mapping relationship between audio and lip movements.
@@
Wang et al(wang2024styletalk++, ) TPAMI’24 ➋➌➏ Unable to delicately control emotions. Using 3DMM as an intermediate variable to convey facial expressions and head movements, and intro -ducing additional reference videos to extract the desired speaking style.
@@
Diffusion DAE-Talker (du2023dae, ) MM’23 High model complexity. It replaces traditional manually crafted intermediate representations with data-driven latent representa -tions obtained from a DAE.
@@
Yu et al(yu2023talking, ) ICCV’23 ➋➒ Poor resolution, high model complexity. Building a corresponding mapping between audio and non-lip representations and training using the diffusion model.
@@
Stypułkowski (stypulkowski2024diffused, ) WACV’24 ➊➑ High model complexity, short video generation duration. The model incorporates motion frame and audio embedding information to capture past movements and future expressions, with an emphasis on the mouth region through an additional lip sync loss.
@@
EmoTalker (zhang2024emotalker, ) ICASSP’24 ➌➑ High model complexity. It achieves emotion-editable talking face generation based on a conditional diffusion model.
@@
VASA-1 (xu2024vasa, ) NIPS’24 High model complexity. Expressive and well-decoupled facial latent space has been constructed, and highly controllable, high -quality generation effects have been achieved based on the Diffusion Transformer.
EmotiveTalk (wang2025emotivetalk, ) CVPR’25 ➌ ➏ High model complexity. Enhances long-duration talking-face generation by introducing a visual-guided audio–expression disen -tanglement strategy and an expressive diffusion backbone, enabling controllable emotional expression.
@@
3D-Model AD-NeRF (guo2021ad, ) ICCV’21 Inadequate control of emotions and latent attributes. The NeRF based approach achieves accurate reproduction of detailed facial components and generates the upper body region.
@@
DFRF (shen2022learning, ) ECCV’22 Lack of emotional and other latent attributes control. Combining audio with 3D perceptual features and proposing an facial deformation module.
@@
AE-NeRF (li2023ae, ) AAAI’24 Lack emotional and other latent attributes control. Facial modeling is divided into NeRF related to audio and unrelated to audio to enhance audio-visual lip synchronization and facial detail.
@@
SyncTalk (peng2023synctalk, ) CVPR’24 Lack controllable emotional intensity regulation. The facial sync controller boosts component coordination, and a portrait generator corrects artifacts, enhancing video details.
@@
Ye et al(ye2024real3d, ) ICLR’24 (zhu2022celebv, ) Lack emotional control and occasional artifacts. Facial and audio information is separately represented using tri-plane, followed by rendering. The gen -erated results are further optimized based on the super-resolution network.
Tang et al(tang2025real, ) IJCV’25 (guo2021ad, )(shen2022learning, ) Lack controllable emotional intensity regulation. This framework enhances reenactment fidelity by decomposing portrait representations into low dimen -sional feature grids, enabling coherent audio-driven head motion and efficient torso modeling.
$\bullet$ Feature Decoupling. The latent feature decoupling and driving methods (bounareli2023hyperreenact, ; bounareli2023stylemask, ; li2023one, ; bounareli2024one, ; yang2023learning, ) aims to disentangle facial features in the latent space of the driving video, replacing or mapping the corresponding latent information to achieve high-fidelity facial reproduction under specific conditions. HyperReenact (bounareli2023hyperreenact, ) uses attribute decoupling, employing a hyper-network to refine source identity features and modify facial poses. StyleMask (bounareli2023stylemask, ) separates facial pose and expression from the identity information of the source image by learning masks and blending corresponding channels in the pre-trained style space S of StyleGAN2. HiDe-NeRF (li2023one, ) employs a deformable neural radiance field to represent a 3D scene, with a lightweight deformation module explicitly decoupling facial pose and expression attributes.
$\bullet$ Self-supervised Learning. Self-supervised learning employs supervisory signals inferred from the intrinsic structure of the data, reducing the reliance on external data labels (tripathy2020icface, ; zeng2020realistic, ; oorloff2023robust, ; zhang2023exploiting, ). Oorloff et al(oorloff2023robust, ) employs self-supervised methods to train an encoder, disentangling identity and facial attribute information of portrait images within the pre-defined latent space itself of a pre-trained StyleGAN2. Zhang et al(zhang2023exploiting, ) utilizes 3DMM to provide geometric guidance, employs pre-computed optical flow to guide motion field estimation, and relies on pre-computed occlusion maps to guide the perception and repair of occluded areas.
3.1.3. Talking Face Generation
@@
In this section, we review current methods from three perspectives: audio/text driven, multimodal conditioned, diffusion-based, and 3D-model Technologies. We also summarize them in Tab. 4.
@@
$\bullet$ Audio/Text Driven. Early methods (fan2015photo, ; chen2018lip, ) perform poorly in terms of generalization and training complexity. After training, the models struggled to generalize to new individuals, requiring extensive conversational data for training new characters. Researchers (chen2019hierarchical, ; prajwal2020lip, ) propose their solutions from various perspectives. However, Most of these methods prioritize generating lip movements aligned with semantic information, overlooking essential aspects like identity and style, such as head pose changes and movement control, which are crucial in natural videos. To address this, MakeItTalk (zhou2020makelttalk, ) decouples input audio information by predicting facial landmarks based on audio and obtaining semantic details on facial expressions and poses from audio signals. SadTalker (zhang2023sadtalker, ) extracts 3D motion coefficients for constructing a 3DMM from audio and uses this to modulate a new 3D perceptual facial rendering for generating head poses in talking videos. Additionally, some methods (wu2023speech2lip, ; wang2023memory, ; lee2024radio, ; fu2023mimic, ) propose their improvement methods, and these will not be detailed one by one. In addition, the emotional expression varies for different texts during a conversation, and vivid emotions are an essential part of real talking face videos (sheng2023stochastic, ; gan2023efficient, ). Recently, some methods (tan2023emmn, ; zhai2023talking, ; gururani2023space, ) extend their previous approaches by incorporating matching between the driving information and corresponding emotions. EMMN (tan2023emmn, ) establishes an organic relationship between emotions and lip movements by extracting emotion embeddings from the audio signal, synthesizing overall facial expressions in talking faces rather than focusing solely on audio for facial expression synthesis. AMIGO (zhai2023talking, ) employs a sequence-to-sequence cross-modal emotion landmark generation network to generate vivid landmarks aided by audio information, ensuring that lips and emotions in the output image sequence are synchronized with the input audio. However, existing methods still lack effective control over the intensity of emotions. In addition, TalkCLIP (ma2025talkclip, ) introduces style parameters, expanding the style categories for text-guided talking video generation. Zhong et al(zhong2023identity, ) propose a two-stage framework, incorporating appearance priors during the generation process to enhance the model’s ability to preserve attributes of the target face. DR2 (zhang2024dr2, ) explores practical strategies for reducing the training workload.
@@
$\bullet$ Multimodal Conditioned. To generate more realistic talking videos, some methods (zhou2021pose, ; liang2022expressive, ; xu2023high, ; wang2023lipformer, ) introduce additional modal information on top of audio-driven methods to guide facial pose and expression. GC-AVT (liang2022expressive, ) generates realistic talking videos by independently controlling head pose, audio information, and facial expressions. This approach introduces an expression source video, providing emotional information during the speech and the pose source video. However, the video quality falls below expectations, and it struggles to handle complex background changes. Xu et al(xu2023high, ) integrate text, image, and audio-emotional modalities into a unified space to complement emotional content in textual information. Multimodal approaches have significantly enhanced the vividness of generated videos, but there is still room for exploration of organically combining information driven by different sources and modalities.
@@
$\bullet$ Diffusion-based. Recently, some methods (mir2023dit, ; du2023dae, ; zhang2024emotalker, ) apply the Diffusion model to the task of talking face generation. For fine-grained talking video generation, DAE-Talker (du2023dae, ) replaces manually crafted intermediate representations, such as facial landmarks and 3DMM coefficients, with data-driven latent representations obtained from a Diffusion Autoencoder (DAE). The image decoder generates video frames based on predicted latent variables. EmoTalker (zhang2024emotalker, ) utilizes a conditional diffusion model for emotion-editable talking face generation. It introduces emotion intensity blocks and the FED dataset to enhance the model’s understanding of complex emotions. Very recently, diffusion models are gaining prominence in talking face generation tasks (stypulkowski2024diffused, ; tian2024emo, ) and video generation tasks (wang2023leo, ; karras2023dreampose, ). Emo (tian2024emo, ) directly predicts video from audio without the need for intermediate 3D components, achieving excellent results. However, the lack of explicit control signals may easily lead to unnecessary artifacts. Based on the Diffusion Transformer architecture, VASA-1 (xu2024vasa, ) finely encodes and reconstructs facial details, constructing an expressive and well-decoupled facial latent space.
@@
$\bullet$ 3D-model Technologies. 3D models, exemplified by NeRF, are gaining traction in talking face generation (guo2021ad, ; shen2022learning, ; li2023ae, ). AD-NeRF (guo2021ad, ) directly feeds features from the input audio signal into a conditional implicit function to generate a dynamic NeRF. AE-NeRF (li2023ae, ) employs a dual NeRF framework to separately model audio-related regions and audio-independent regions. Furthermore, some methods (peng2023synctalk, ; ye2024real3d, ) adopt Tri-Plane to represent facial and audio attributes. Synctalk (peng2023synctalk, ) models and renders head motion using a tri-plane hash representation, and then further enhances the output quality using a portrait synchronization generator. Very recently, 3D Gaussian Splatting (kerbl20233d, ) also been widely applied to this task.
@@
Table 5. Overview of facial attribute editing methods. Notations: ➊ FFHQ, ➋ CelebA, ➌ CelebA-HQ, ➍ CelebAMask-HQ, ➎ VoxCeleb, ➏ CelebAText-HQ, ➐ LFW, ➑ $M$ CelebA-HQ, ➒ CARLA, ➓ Multi-PIE. In addition, abbreviations are used in the table: SIGGRAPH (SIG.), GANs (G.), Diffusion (D.), Transformer (T.).
@@
GeneGAN (zhou2017genegan, ) BMVC’17 G. ➋➓ In early attribute editing, separate models were trained for specific attributes. The key idea was to reassemble attribute vectors in the latent space, achieving successful editing.
@@
SC-FEGAN (jo2019sc, ) ICCV’19 G. Users can generate high-quality edited output images by freely sketching parts of the source image.
@@
AttGAN (he2019attgan, ) TIP’19 G. Applying attribute classification constraints to generated images has validated the drawbacks of enforcing stringent attribute independence constraints in latent representations.
@@
Shen et al(shen2020interpreting, ) CVPR’20 G. Thoroughly investigated how to encode different semantics in the latent space and explored the disentanglement between various semantics to achieve precise control over facial attributes.